if n is a positive integer , let S(n) denote the sum of the positive divisors of n, including n, G(n) is the greatest divisor of n. If H(n)=(G(n))/(S(n)) then which of the following is the largest H(2010) or H(2011)

Grade:12

1 Answers

Sher Mohammad IIT Delhi
9 years ago
S(2010)=1+2+3+67=73
S(2011)=1
G(2011)=1
G(2010)=67
H(2010)=67/73
H(2011)=1
H(2011) is greater.